A particular shade of paint is made by mixing 5 parts red paint with 7 parts blue paint. To make this shade, Shannon mixed 12 quarts of blue pain with 8 quarts of red paint. Did Shannon mix the correct shade? Explain.
Answer by jim_thompson5910(35256)

If Shannon mixed the correct shade, then the ratios would be equal. Let's test that

(7 parts blue)/(5 parts red) = (12 parts blue)/(8 parts red)

7/5 = 12/8

7*8 = 5*12 ... cross multiply

56 = 60

Since the last equation is false, this means the first equation is false. So the two ratios are NOT the same. That means that the two shades will be different. She did not mix the correct shade.